OpenStudy (anonymous):

Write an equation in slope-intercept form for the line described. Passes through (-8,-2), m= 5/2

OpenStudy (ranga):

Slope intercept form is: y = mx + b where m is the slope and b is the y-intercept. Put the m value above. Since the line passes through (-8,-2), put x = -8 and y = -2 and solve for b.

OpenStudy (anonymous):

y= 5/2 + 6 ?

OpenStudy (ranga):

y = 5/2x + b when x = -8, y = -2 -2 = 5/2(-8) + b solve for b.

OpenStudy (anonymous):

b= -40 ?

OpenStudy (ranga):

no.

OpenStudy (ranga):

-2 = 5/2(-8) + b -2 = -40/2 + b -2 = -20 + b -2+20 = b b = 18

OpenStudy (anonymous):

I see thanks!

OpenStudy (ranga):

yw.